OpenStudy (anonymous):

6(4/5)n-(8/9)=(7/15)n

Directrix (directrix):

6(4/5) n - (8/9 )= (7/15) n Multiply both sides by 45 ...

Directrix (directrix):

45[6(4/5) n] - 45(8/9 )=45 (7/15) n 216 n - 40 = 21 n ----> Please check

Directrix (directrix):

216 n - 21 n = 40

Directrix (directrix):

195 n = 40 n = 40/ 195 = 8 / 39

Directrix (directrix):

Question?